# Derivative-Aligned Anticipation of Forbush Decreases from Entropy and Fractal Markers

### v1

@@ 6 Conclusions
We introduced and validated an invariant–based, derivative–aligned framework for anticipating FDs in neutron–monitor records. Sliding–window features are converted into station–level marker series and analyzed relative to the time of the minimum of the smoothed count derivative, yielding leads in minutes and an operational coverage metric based on a robust zz–score with bilateral threshold and persistence on pre–onset windows. Across three events and multiple stations, a compact panel—Hurst (via DFA), Katz fractal dimension, Shannon and sample entropy, and Lempel–Ziv complexity—consistently precedes FD onsets with substantial station coverage, offering practically useful lead times under both extreme and rapid morphologies. Event–dependent behavior of permutation and spectral entropy underscores the value of ensembles rather than single–feature triggers.
@@
Minute–resolution neutron–monitor data were obtained from the Neutron Monitor Database (NMDB, https://www.nmdb.eu). In accordance with NMDB terms of use, we do not redistribute raw counts; the same station series for the three events (2023-04-23, 2024-03-24, 2024-05-10) can be retrieved directly from NMDB.
All analysis code to reproduce the figures, tables, and statistics is openly available at https://github.com/dosquisd/NMDB/tree/main. For exact reproducibility, we recommend citing a tagged release or commit hash of the repository and recording the NMDB query metadata used to export the minute–resolution series.

### v2

@@ 6 Conclusions
We presented and evaluated an invariant-based, derivative-aligned framework to anticipate Forbush decreases (FDs) in neutron-monitor count-rate records using a fully automated, station-wise procedure. The method converts EWM-smoothed count series into sliding-window invariants and extracts event-relative timing by aligning each station to an operational reference time $t_{0}$, defined as the minimum of the EWM-smoothed first difference of the count rate. Marker leads are then quantified in minutes from sustained excursions in the derivatives of the invariant time series within a pre-onset search window, and summarized through detection coverage and robust distributional statistics (medians and IQRs) across the station network.
Applied to two one-minute-cadence FD episodes (2023-04-23 and 2024-05-10) across 28 stations per event, the results show that a compact CORE panel spanning information-theoretic measures (Shannon, spectral, sample, and approximate entropy), geometric descriptors (Katz and Higuchi fractal dimension), and complexity/embedding indicators (Lempel–Ziv and correlation dimension) provides reproducible anticipatory signatures relative to $t_{0}$. Across the network, these invariants achieve high detection coverage (typically $\sim$78–93%) and predominantly negative station-wise lags, with typical median lead times on the order of several hours. While lead distributions remain heterogeneous across stations, the combination of high coverage and consistent negative medians supports the interpretation that anticipatory structure is a network-level signal rather than a single-station artifact.
Finally, the onset diagnostic and the marker time-series examples demonstrate that the reported leads correspond to identifiable, persistent features in the invariant trajectories rather than isolated extrema. Together, the onset diagnostics, pooled/event-wise rankings, and station-wise distributions support the CORE panel as a parsimonious basis for early flagging in a multi-station setting. Future work should extend the evaluation to larger event sets spanning solar-cycle phases and rigidity bands, and explore adaptive/multiscale windows and multimodal fusion with upstream solar-wind and IMF context to improve calibration and operational alerting.
@@
Minute–resolution neutron–monitor data were obtained from the Neutron Monitor Database (NMDB, https://www.nmdb.eu). In accordance with NMDB terms of use, we do not redistribute raw counts; the same station series for the Two events (2023-04-23 and 2024-05-10) can be retrieved directly from NMDB. All analysis code to reproduce the figures, tables, and statistics is openly available at https://github.com/sierraporta/FD-characterization-topological-approach/tree/main. For exact reproducibility, we recommend citing a tagged release or commit hash of the repository and recording the NMDB query metadata used to export the minute–resolution series.